prettygirl15:

The ratio of green marbles to yellow marbles is 24 green marbles to 9 yellow marbles. If there are 3 yellow marbles, how many green marbles are there?

prettygirl15:

ShoaibChaudhry:

Ok Again Wat is 9 dived by 3?

ShoaibChaudhry:

So bacily we go from 9 to 3 right And ration are always porportional

ShoaibChaudhry:

And then 24 divied by 3

ShoaibChaudhry:

ther for ur answer is 8

jhonyy9:

using the rule of 3 simple you can write 24 green to 9 yellow x green to 3 yellow cross multiplie and calcule the x value 9x = 24*3 x = 24*3/9 x = ?

ShoaibChaudhry:

Buhh Wouldnt that be A bit to hard to understand

Lillithluv:

i am confused on that aswell but i was going to say divide 9 and 3